A random sample of 16 values of x, obtained from a N(u, Q2)distribution, had a sample mean of 4:2 and a sample variance of 3: Find the 95% confidence interval for u



1
Expert's answer
2021-07-06T15:27:52-0400

The critical value for "\alpha=0.05" and "df=n-1=16-1=15" degrees of freedom is "t_c=2.131449."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(4.2-2.131449\times\dfrac{
\sqrt{3}}{\sqrt{16}}, 4.2+2.131449\times\dfrac{\sqrt{3}}{\sqrt{16}})"

"=(3.277, 5.123)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "3.277<u<5.123," which indicates that we are 95% confident that the true population mean "u" is contained by the interval "(3.277, 5.123)."